OpenStudy (anonymous):

how do you solve -12-x=-16

OpenStudy (anonymous):

-12-x=-16 -x-12=-16 x=4

OpenStudy (anonymous):

take the 12 and add it to both sides. now you have -x equals -4 . From here divide -x on both sides and two negatives equal a positive so now you get x equals a positive 4.

OpenStudy (anonymous):